# A Measurement Study of Model Context Protocol

### v1

@@ Abstract.
Model Context Protocol, MCP, Crawler, Measurement
copyright: rightsretainedjournal: POMACSjournalyear: 2021journalvolume: 5journalnumber: 2article: 14publicationmonth: 6doi: 10.1145/3460081ccs: Network performance evaluation Network Measurementccs: Security and privacy Systems security Operating systems security Mobile platform securityccs: Security and privacy Software and application security Software security engineeringccs: Software and its engineering Software system structures Real-time systems software
@@ 1. Introduction
The recent surge of large language model (LLM)–powered applications (openai2023plugins, ; anthropic2024claude, ) has brought new demands for interoperability, modularity, and extensibility. The Model Context Protocol (MCP) has emerged as a promising standard to meet these needs, offering a lightweight interface for connecting clients, servers, and external resources  (openai2024mcp, ). MCP specifies a lightweight yet expressive mechanism for describing, discovering, and invoking external capabilities within a model’s operational context. Its ambition is to become for LLM integration what HTTP was for the Web (fielding2000architectural, ) or what USB became for peripheral devices (platt1999plugplay, ), a unifying protocol layer that ensures portability and reusability across diverse platforms. The significance of such a standard has also been recognized in broader AI governance contexts: international standardization bodies such as ISO/IEC JTC 1/SC 42 have emphasized the need for interoperability and sustainable integration frameworks for AI technologies (isoiec2025sc42, ).
However, despite its promise, the trajectory of the MCP ecosystem remains highly uncertain. Marketplaces may list thousands of entries, but it is unclear how many represent actively maintained, production-quality projects. For instance, users on Reddit have questioned whether MCP is simply another overhyped standard destined to fade once the novelty wears off, noting that “most of MCPs, like sequential thinking, don’t really need to be MCP and are not a good fit” (reddit2025mcp, ). Without a systematic perspective, it is difficult to determine whether MCP is undergoing genuine, sustainable growth or merely experiencing a transient surge driven by novelty and hype. At the same time, MCP servers frequently integrate with sensitive resources such as authentication systems, proprietary APIs, and personal data connectors. While such deployments could enable secure interoperability, they also risk exposing users to privacy and security vulnerabilities if not carefully safeguarded. MCP clients, which act as the bridge between servers and LLMs, further influence the ecosystem’s trajectory through their choice of communication protocols and connection modes. Some developers advocate for SSE as the natural default for future interoperability, while others favor the simplicity of stdio for debugging and lightweight use.
@@
  • (Server – Security and Privacy Posture). Our analysis of 8,060 valid MCP servers highlights multiple structural risks. On dependencies, we found strong monocultures: e.g., Java servers overwhelmingly use Spring (spring-boot, spring-core, spring-web), meaning vulnerabilities such as SpringShell could cascade widely. Python and TypeScript servers rely on pydantic and zod, which improve input safety, but Go and Rust servers lack equivalent safeguards, relying on manual validation. Maintenance also varies: 40.9% of servers were updated within 90 days, but 21.9% had been inactive for over a year, creating a long tail of unpatched projects. Functionally, 11.2% of servers expose sensitive APIs, with authentication-related services making up 43% of this group. These are better maintained (58% active in 90 days), but their exposure amplifies the consequences of misconfiguration or compromise. Together, these findings show that while good practices exist, the ecosystem remains highly vulnerable to supply-chain attacks, abandonment risks, and privacy exposures.
@@
  • First large-scale empirical study. We present the first large-scale empirical study of the MCP ecosystem, covering six major markets and systematically analyzing 8,060 servers and 341 clients. The collected dataset has been made publicly available at https://github.com/zhuaiballl/mcp_collection.
  • Measurement framework and dataset. We design and implement MCPCrawler, a dedicated measurement framework for discovering, filtering, and analyzing MCP projects, and release the resulting dataset to support reproducible research and future studies. MCPCrawler is open-sourced at https://github.com/zhuaiballl/mcpc.
@@ 2.1. MCP Overview
The Model Context Protocol (MCP) is an emerging open standard for connecting large language model (LLM) applications to external data sources, tools, and services (anthropic2024mcp, ; tajwar2024preferencefinetuningllmsleverage, ; openaitoolcalling, ). It serves as a “standardization layer” that lets an LLM interact with live data and APIs much like a USB-C port links a computer to peripherals (wei2022emergent, ; cai2024llmtool, ). MCP treats the AI application as an MCP Host that can “mount” external MCP servers via one or more MCP clients (glama2024doc, ; smithery2024market, ). Each MCP server exposes contextual resources (data), prompts (templates), and tools (functions) to the connected AI system, enabling the model to retrieve fresh information and invoke actions securely (qin2024survey, ; iqbal2024llmplatformsecurityapplying, ; bommasani2021opportunities, ). For example, an MCP server might expose a “weather/current” tool or a “database/schema” resource; the AI application’s MCP client discovers these offerings and incorporates them into the model’s toolset (yao2023react, ; schick2023toolformer, ).
@@
MCP Architecture. MCP is designed around a client–server paradigm that leverages stateful JSON-RPC 2.0 communication (jsonrpc, ; fielding2000architectural, ). Rather than being a monolithic framework, MCP provides a structured way for LLM-equipped applications to discover, access, and coordinate external resources. This design reflects the broader trend of treating LLMs not only as text generators but also as orchestrators that interact with heterogeneous tools and services. As shown in Figure 1, key components of the architecture include:
@@
  • MCP Host (Application): The LLM-equipped application (e.g., Claude Desktop, a chat interface, or an IDE plugin) that initiates connections to servers. The host creates a separate MCP client for each server it connects to (anthropic2024mcp, ; glama2024doc, ).
@@
  • MCP Client: A client object embedded in the host, responsible for a one-to-one connection with an MCP server. A host connecting to $n$ servers will have $n$ MCP client instances (openaitoolcalling, ).
@@
  • MCP Server: A service (local or remote) that provides contextual data and executable tools (iqbal2024llmplatformsecurityapplying, ; tajwar2024preferencefinetuningllmsleverage, ). Servers can run locally using stdio transport or on cloud infrastructure via streamable HTTP (jsonrpc, ).
@@
  • (Optional) Context Resolver: Some advanced agent frameworks introduce a higher-level resolver that routes queries to the most relevant server based on the domain of the request (liu2023agentbench, ; guo2024llmbased, ). For example, if a user asks “summarize the latest customer support tickets”, the resolver can direct the request to a server connected to the company’s ticketing database (e.g., Zendesk API). Conversely, a request such as “generate a UML diagram from this code snippet” would be dispatched to a code-analysis server or IDE plugin. By abstracting this routing logic, the resolver allows the host to handle heterogeneous queries seamlessly without hardcoding which server should be called.
MCP Workflow. The MCP workflow captures the end-to-end interaction pattern by which an LLM, via its client, establishes connections, discovers tools, and invokes server functions to fulfill user requests. The LLM, acting on the user’s natural language instruction, delegates execution to the MCP client, which initializes the connection, exchanges protocol versions and capability sets with the server, and subscribes to notifications (glama2024doc, ; anthropic2024mcp, ). After initialization, the client performs tool and resource discovery (e.g., via $\mathsf{tools/list}$ or $\mathsf{resources/list}$ requests) (iqbal2024llmplatformsecurityapplying, ; yao2023react, ). The host application aggregates these listings into a unified tool registry that the LLM can query when deciding how to fulfill user requests. When the user issues a task (e.g., “find recent papers” or “summarize a document”), the LLM selects a tool and triggers its execution by sending a $\mathsf{tools/call}$ request with the tool’s URI and arguments. The server executes the corresponding function (such as querying a database or performing a web search) and returns a structured response (schick2023toolformer, ; cai2024llmtool, ), which is then processed by the LLM and surfaced back to the user in natural language. In addition, servers can proactively push real-time notifications when their available primitives change, ensuring that both the LLM and user always operate over the latest tool state (jsonrpc, ).
@@
Deployment and Implementation of MCP. MCP relies on two transport layers to accommodate different deployment settings: for local servers, stdio offers low-latency communication without network overhead, while for remote servers, HTTP/S with optional Server-Sent Events supports streaming and standard authentication (anthropic2024mcp, ; glama2024doc, ). Beyond transport, MCP is inherently stateful: once initialized, the client and server maintain a session that allows multiple RPC calls to be issued over the same channel (tajwar2024preferencefinetuningllmsleverage, ). This design enables AI applications to “carry their context” across tools and environments, ensuring interoperability in heterogeneous ecosystems (cai2024llmtool, ; hassouna2024llmagentumf, ). Building on this foundation, different implementations have emerged that extend MCP’s core. Some projects enrich the protocol into agent frameworks with advanced features such as namespace resolution or decentralized data networks (liu2023agentbench, ; guo2024llmbased, ), while others embed reasoning logic, delegating sub-questions back to the model for refinement (yao2023react, ; schick2023toolformer, ). Despite such variations, all implementations adhere to the same underlying handshake, discovery, execution, and notification flows that define MCP.
@@ 2.2. MCP Markets
A central part of the MCP ecosystem is the emergence of dedicated marketplaces (sometimes referred to as registries), which act as catalogs for MCP-compatible servers and clients. These platforms provide indexing, categorization, and in some cases direct hosting of connectors (smithery2024market, ; glama2024doc, ; pulsemcp2024, ). By doing so, they reduce duplication of effort and enable developers or agent frameworks to reuse existing connectors rather than building new ones from scratch—mirroring the dynamics seen in plugin ecosystems for web browsers, IDEs, and LLM agents (liu2023agentbench, ; li2024survey, ). Well-known examples include Smithery, MCP.so, Glama.ai, PulseMCP, MCP Market, and Cursor Directory (smithery2024market, ; glama2024doc, ; pulsemcp2024, ).
MCP marketplaces adopt different deployment models, reflecting a trade-off between local control and cloud-managed convenience. Some markets support local mode, where server code is fetched and run in the user’s own environment, giving users full control of credentials and runtime (smithery2024market, ; glama2024doc, ). Others provide hosted mode, offering cloud-hosted MCP servers with managed runtime and stable endpoints, which reduce setup cost but delegate execution control to the hosting provider (glama2024doc, ; pulsemcp2024, ). This design resembles trends in serverless computing and API marketplaces (buyya2008market, ; javed201652, ). Other registries, such as MCP.so or Cursor Directory, primarily focus on discovery and indexing rather than deployment (pulsemcp2024, ). Table 1 summarizes the major MCP marketplaces, comparing their API/SDK support, deployment options, protocol coverage, and scale of discoverability (smithery2024market, ; glama2024doc, ). These marketplaces collectively illustrate how MCP is evolving into a multi-platform ecosystem similar to established software markets (buyya2008market, ; li2024survey, ).
@@ 4.1. Challenges and Insights
C1. Market-level: Data Heterogeneity and Access Restrictions. MCP servers and MCP clients are indexed by multiple markets (e.g., Glama.ai, MCP.so, GitHub community repositories, PulseMCP). These markets expose project metadata through heterogeneous mechanisms such as JSON APIs, HTML listings, or keyword-based search results. Their metadata schemas also diverge: some markets emphasize quality ratings (e.g., security and license tiers), while others omit such fields entirely. In addition, markets impose strict access constraints, including hard result caps, API rate limits, or Cloudflare-based CAPTCHA challenges. Such heterogeneity and restrictions complicate normalized aggregation, create coverage gaps, and reduce reproducibility of market-level analyses.
@@ 4.2.2. Server Resolver
GitHub Repositories Unstructured README files Search keyword bias Automatic keyword expansion
@@ 4.3. Performance of MCPCrawler
  • Data Quality and Noise Filtering. Out of the 16,950 MCP server entries collected, 8,635 were identified as low-value or invalid (e.g., placeholder repositories, inactive forks). Meanwhile, 343 out of 363 MCP client entries collected were identified as valid. This finding suggests that MCP clients tend to be more consistently maintained and of higher quality compared to MCP servers. MCPCrawler’s noise filtering module automatically excluded 50.9% of these entries. Manual validation of a 500-entry sample confirmed 93.5% accuracy in invalid server detection.
@@ 5. Measurement of MCP Markets
In this section, we conduct a measurement study of the major markets that index and distribute MCP-related projects. Platforms such as Glama.ai, MCP.so, PulseMCP, and GitHub repositories act as the primary gateways for MCP servers and clients, but they differ widely in scope, metadata practices, and access restrictions. These differences directly shape what projects become visible to the ecosystem and how representative each market truly is. By systematically quantifying market size, examining overlaps across sources, and characterizing the diversity of listed entries, our analysis directly addresses RQ1 on the overall scale and distribution of the MCP ecosystem.
@@ 6.1. Measurement of Library Ecosystems for Supply-Chain Security Assessment
To answer RQ2, we next examine the libraries that MCP servers depend on, with the goal of understanding not only their engineering practices but also their security posture. Libraries are a critical dimension of risk: they may introduce vulnerabilities through unsafe defaults, amplify supply-chain exposure, or conversely, provide safeguards such as schema validation and secure communication. Measuring which libraries are most widely used therefore helps reveal where sensitive functionality is concentrated and whether best practices are being adopted. Figure 6 presents the top 20 libraries used by MCP servers in each programming language. To produce this figure, we extracted declared dependencies from server repositories, normalized package names across ecosystems, and computed their frequency. We then visualized the results as bar charts grouped by language, where each bar corresponds to the number of servers importing a given library. This representation makes it straightforward to compare the prevalence of libraries both within and across language stacks.

### v3

@@ Abstract.
Model Context Protocol, MCP, MCP Dataset, Crawler, Measurement
copyright: rightsretainedccs: Network performance evaluation Network Measurementccs: Security and privacy Systems security Operating systems security Mobile platform securityccs: Security and privacy Software and application security Software security engineeringccs: Software and its engineering Software system structures Real-time systems software
@@ 1. Introduction
The recent surge of large language model (LLM)–powered applications (openai2023plugins; anthropic2024claude) has brought new demands for interoperability, modularity, and extensibility. The Model Context Protocol (MCP) has emerged as a promising standard to meet these needs, offering a lightweight interface for connecting clients, servers, and external resources  (introducingmcp; li2025largelanguagemodelstrusted; li2025urgentlyneedprivilegemanagement). MCP specifies a lightweight yet expressive mechanism for describing, discovering, and invoking external capabilities within a model’s operational context. Its ambition is to become for LLM integration what HTTP was for the Web (fielding2000architectural) or what USB became for peripheral devices (windowsplugplay), a unifying protocol layer that ensures portability and reusability across diverse platforms. The significance of such a standard has also been recognized in broader AI governance contexts: international standardization bodies such as ISO/IEC JTC 1/SC 42 have emphasized the need for interoperability and sustainable integration frameworks for AI technologies (isoiec2025sc42).
However, despite its promise, the trajectory of the MCP ecosystem remains highly uncertain. Marketplaces may list thousands of entries, but it is unclear how many represent actively maintained, production-quality projects. For instance, users on Reddit have questioned whether MCP is simply another overhyped standard destined to fade once the novelty wears off, noting that “most of MCPs, like sequential thinking, don’t really need to be MCP and are not a good fit” (reddit2025mcp). Without a systematic perspective, it is difficult to determine whether MCP is undergoing genuine, sustainable growth or merely experiencing a transient surge driven by novelty and hype. At the same time, MCP servers frequently integrate with sensitive resources such as authentication systems, proprietary APIs, and personal data connectors. While such deployments could enable secure interoperability, they also risk exposing users to privacy and security vulnerabilities if not carefully safeguarded. MCP clients, which act as the bridge between servers and LLMs, further influence the ecosystem’s trajectory through their choice of communication protocols and connection modes. Some developers advocate for SSE as the natural default for future interoperability, while others favor the simplicity of stdio for debugging and lightweight use.
@@
  • (Server – Security and Privacy Posture). Our analysis of 8,060 valid MCP servers highlights multiple structural risks. On dependencies, we found strong monocultures: e.g., Java servers overwhelmingly use Spring (spring-boot, spring-core, spring-web), meaning vulnerabilities such as SpringShell could cascade widely. Python and TypeScript servers rely on pydantic and zod, which improve input safety, but Go and Rust servers lack equivalent safeguards, relying on manual validation. Maintenance also varies: 40.9% of servers were updated within 90 days, but 21.9% had been inactive for over a year, creating a long tail of unpatched projects. Functionally, 11.2% of servers contain code that potentially invokes sensitive APIs, with authentication-related services comprising 43% of this group. Their inclusion of high-impact APIs increases the potential consequences of misconfiguration or compromise. Together, these findings show that while good practices exist, the ecosystem remains highly vulnerable to supply-chain attacks, abandonment risks, and privacy exposures.
@@
  • First large-scale empirical study. We present the first large-scale empirical study of the MCP ecosystem, covering six major markets and systematically analyzing 8,060 servers and 341 clients.
  • Measurement framework and dataset. We design and implement MCPCrawler, a dedicated measurement framework for discovering, filtering, and analyzing MCP projects, and release the resulting dataset to support reproducible research and future studies.
@@
Dataset and Tool. MCPCrawler is open-sourced at https://github.com/zhuaiballl/mcpc. The collected dataset has been made publicly available at https://github.com/zhuaiballl/mcp_collection.
@@ 2.1. MCP Overview
The Model Context Protocol (MCP) is an emerging open standard for connecting large language model (LLM) applications to external data sources, tools, and services (anthropic2024mcp; tajwar2024preferencefinetuningllmsleverage; openaitoolcalling; surveymcp). It serves as a “standardization layer” that lets an LLM interact with live data and APIs much like a USB-C port links a computer to peripherals (wei2022emergent; cai2024llmtool). MCP treats the AI application as an MCP Host that can “mount” external MCP servers via one or more MCP clients (glama2024doc; smithery2024market). Each MCP server exposes contextual resources (data), prompts (templates), and tools (functions) to the connected AI system, enabling the model to retrieve fresh information and invoke actions securely (qin2024survey; iqbal2024llmplatformsecurityapplying; bommasani2021opportunities). For example, an MCP server might expose a “weather/current” tool or a “database/schema” resource; the AI application’s MCP client discovers these offerings and incorporates them into the model’s toolset (yao2023react; schick2023toolformer).
@@
MCP Architecture. MCP is designed around a client–server paradigm that leverages stateful JSON-RPC 2.0 communication (jsonrpc; fielding2000architectural). Rather than being a monolithic framework, MCP provides a structured way for LLM-equipped applications to discover, access, and coordinate external resources. This design reflects the broader trend of treating LLMs not only as text generators but also as orchestrators that interact with heterogeneous tools and services. As shown in Figure 1, key components of the architecture include:
@@
  • MCP Host (Application): The LLM-equipped application (e.g., Claude Desktop, a chat interface, or an IDE plugin) that initiates connections to servers. The host creates a separate MCP client for each server it connects to (anthropic2024mcp; glama2024doc).
@@
  • MCP Client: A client object embedded in the host, responsible for a one-to-one connection with an MCP server. A host connecting to $n$ servers will have $n$ MCP client instances (openaitoolcalling).
@@
  • MCP Server: A service (local or remote) that provides contextual data and executable tools (iqbal2024llmplatformsecurityapplying; tajwar2024preferencefinetuningllmsleverage). Servers can run locally using stdio transport or on cloud infrastructure via streamable HTTP (jsonrpc).
@@
  • (Optional) Context Resolver: Some advanced agent frameworks introduce a higher-level resolver that routes queries to the most relevant server based on the domain of the request (liu2023agentbench; guo2024llmbased). For example, if a user asks “summarize the latest customer support tickets”, the resolver can direct the request to a server connected to the company’s ticketing database (e.g., Zendesk API). Conversely, a request such as “generate a UML diagram from this code snippet” would be dispatched to a code-analysis server or IDE plugin. By abstracting this routing logic, the resolver allows the host to handle heterogeneous queries seamlessly without hardcoding which server should be called.
@@
MCP Workflow. The MCP workflow captures the end-to-end interaction pattern by which an LLM, via its client, establishes connections, discovers tools, and invokes server functions to fulfill user requests. The LLM, acting on the user’s natural language instruction, delegates execution to the MCP client, which initializes the connection, exchanges protocol versions and capability sets with the server, and subscribes to notifications (glama2024doc; anthropic2024mcp). After initialization, the client performs tool and resource discovery (e.g., via $\mathsf{tools/list}$ or $\mathsf{resources/list}$ requests) (iqbal2024llmplatformsecurityapplying; yao2023react). The host application aggregates these listings into a unified tool registry that the LLM can query when deciding how to fulfill user requests. When the user issues a task (e.g., “find recent papers” or “summarize a document”), the LLM selects a tool and triggers its execution by sending a $\mathsf{tools/call}$ request with the tool’s URI and arguments. The server executes the corresponding function (such as querying a database or performing a web search) and returns a structured response (schick2023toolformer; cai2024llmtool), which is then processed by the LLM and surfaced back to the user in natural language. In addition, servers can proactively push real-time notifications when their available primitives change, ensuring that both the LLM and user always operate over the latest tool state (jsonrpc).
Deployment and Implementation of MCP. MCP relies on two transport layers to accommodate different deployment settings: for local servers, stdio offers low-latency communication without network overhead, while for remote servers, HTTP/S with optional Server-Sent Events supports streaming and standard authentication (anthropic2024mcp; glama2024doc). Beyond transport, MCP is inherently stateful: once initialized, the client and server maintain a session that allows multiple RPC calls to be issued over the same channel (tajwar2024preferencefinetuningllmsleverage). This design enables AI applications to “carry their context” across tools and environments, ensuring interoperability in heterogeneous ecosystems (cai2024llmtool; hassouna2024llmagentumf). Building on this foundation, different implementations have emerged that extend MCP’s core. Some projects enrich the protocol into agent frameworks with advanced features such as namespace resolution or decentralized data networks (liu2023agentbench; guo2024llmbased), while others embed reasoning logic, delegating sub-questions back to the model for refinement (yao2023react; schick2023toolformer). Despite such variations, all implementations adhere to the same underlying handshake, discovery, execution, and notification flows that define MCP.
@@ 2.2. MCP Markets
A central part of the MCP ecosystem is the emergence of dedicated marketplaces (sometimes referred to as registries), which act as catalogs for MCP-compatible servers and clients. These platforms provide indexing, categorization, and in some cases direct hosting of connectors (smithery2024market; glama2024doc; pulsemcp2024). By doing so, they reduce duplication of effort and enable developers or agent frameworks to reuse existing connectors rather than building new ones from scratch—mirroring the dynamics seen in plugin ecosystems for web browsers, IDEs, and LLM agents (liu2023agentbench; li2024survey). Well-known examples include Smithery, MCP.so, Glama.ai, PulseMCP, MCP Market, and Cursor Directory (smithery2024market; glama2024doc; pulsemcp2024).
MCP marketplaces adopt different deployment models, reflecting a trade-off between local control and cloud-managed convenience. Some markets support local mode, where server code is fetched and run in the user’s own environment, giving users full control of credentials and runtime (smithery2024market; glama2024doc). Others provide hosted mode, offering cloud-hosted MCP servers with managed runtime and stable endpoints, which reduce setup cost but delegate execution control to the hosting provider (glama2024doc; pulsemcp2024). This design resembles trends in serverless computing and API marketplaces (buyya2008market; javed201652). Other registries, such as MCP.so or Cursor Directory, primarily focus on discovery and indexing rather than deployment (pulsemcp2024). Table 1 summarizes the major MCP marketplaces, comparing their API/SDK support, deployment options, protocol coverage, and scale of discoverability (smithery2024market; glama2024doc). These marketplaces collectively illustrate how MCP is evolving into a multi-platform ecosystem similar to established software markets (buyya2008market; li2024survey).
@@ 4.1. Challenges and Insights
C1. Market-level: Data Heterogeneity and Access Restrictions. MCP servers and MCP clients are indexed by multiple markets (e.g., Glama.ai, MCP.so, PulseMCP). These markets expose project metadata through heterogeneous mechanisms such as JSON APIs, HTML listings, or keyword-based search results. Their metadata schemas also diverge: some markets emphasize quality ratings (e.g., security and license tiers), while others omit such fields entirely. In addition, markets impose strict access constraints, including hard result caps, API rate limits, or Cloudflare-based CAPTCHA challenges. Such heterogeneity and restrictions complicate normalized aggregation, create coverage gaps, and reduce reproducibility of market-level analyses.
@@ 4.3. Performance of MCPCrawler
  • Data Quality and Noise Filtering. Out of the 16,950 MCP server entries collected, 8,635 were identified as low-value or invalid (e.g., placeholder repositories, inactive forks). Meanwhile, 341 out of 363 MCP client entries collected were identified as valid. This finding suggests that MCP clients tend to be more consistently maintained and of higher quality compared to MCP servers. MCPCrawler’s noise filtering module automatically excluded 50.9% of these entries. Manual validation of a 500-entry sample confirmed 93.5% accuracy in invalid server detection.
@@ 5. Measurement of MCP Markets
In this section, we conduct a measurement study of the major markets that index and distribute MCP-related projects. Platforms such as Glama.ai, MCP.so, and PulseMCP act as the primary gateways for MCP servers and clients, but they differ widely in scope, metadata practices, and access restrictions. These differences directly shape what projects become visible to the ecosystem and how representative each market truly is (extension). By systematically quantifying market size, examining overlaps across sources, and characterizing the diversity of listed entries, our analysis directly addresses RQ1 on the overall scale and distribution of the MCP ecosystem.
@@ 6.1. Measurement of Library Ecosystems for Supply-Chain Security Assessment
To answer RQ2, we next examine the libraries that MCP servers depend on, with the goal of understanding not only their engineering practices but also their security posture. Libraries are a critical dimension of risk: they may introduce vulnerabilities through unsafe defaults, amplify supply-chain exposure, or conversely, provide safeguards such as schema validation and secure communication(smallworld). Measuring which libraries are most widely used therefore helps reveal where sensitive functionality is concentrated and whether best practices are being adopted. Figure 6 presents the top 20 libraries used by MCP servers in each programming language. To produce this figure, we extracted declared dependencies from server repositories, normalized package names across ecosystems, and computed their frequency. We then visualized the results as bar charts grouped by language, where each bar corresponds to the number of servers importing a given library. This representation makes it straightforward to compare the prevalence of libraries both within and across language stacks.